I'm lost on part b, please show all work/steps & write clearly please thanks!
Answer #1

a)
Power output, P_out = 6.5 kW

= 6500 J/s

Workdone per every second, W = 6500 J

efficienicy, e = 27 % = 0.27

let Qin is the amount energy delivred to the engine per second.

use, e = W/Qin

Qin = W/e

= 6500/0.27

= 24074 J <<<<<<<<<<----------------Answer

b) amount of energy ejected per second, Q_out = Qin - W

= 24074 - 6500

= 17574 J

no of cycles per second, N = Amount of energy ejected per second/amount of energy ejected per cycle

= 17574/3400

N = 5.2 cycles per second <<<<<<<<<<----------------Answer